OpenStudy (anonymous):

heidi left a 25% tip of $11.25 on a meal. what was the cost of the meal before the tip? what was the total cost of the meal including the tip?

OpenStudy (lukecrayonz):

$45

OpenStudy (anonymous):

huh?howd you get that, and theres TWO parts to it still, thanks

zepdrix (zepdrix):

cost of the meal, let that equal x .25x = 11.25 <-- this is the size of the tip, 25% of x, (25% of the price of the meal was 11.25). (1/4)x= 11.25 x=(4)*11.25=45 Make sense? :)

OpenStudy (anonymous):

yes, thanks!!!:)

OpenStudy (anonymous):

got it, 56.25 with the tip, 45.00 without

zepdrix (zepdrix):

yay

OpenStudy (lukecrayonz):

Figured you could do the 2nd part by yourself lol